OpenStudy (anonymous):

3 over 10 x = 13 - x How does this equal 10 ?

OpenStudy (anonymous):

What do you mean 3 over 10x? You mean 10x/3=13-x ? can you rephrase it?

OpenStudy (anonymous):

3 tenths x = 13 - x ?

OpenStudy (anonymous):

im working it on paper one second.

OpenStudy (anonymous):

10 is the answer because that's what my graphing calculator gave me. I'm trying to see how the answer became 10.

OpenStudy (anonymous):

So using cross products 3x = 10 (13 -x) 3x = 130 - 10x so 13x = 130 x = 10

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

wow good job, i did it wrong :P